# Robustly optimal dynamics for active matter reservoir computing

### v2

@@ III.2 Signatures from microscopic dynamics with and without driving
Figure 3: Characterizing intrinsic microscopic relaxation dynamics (a,b) as well as dynamical correlations under driving (c,d). Symbols correspond to parameter combinations changing speed-controller settings $(K_{sc},s)$ (and resulting damping dynamics) as shown in Fig. 2. (a,b) Intrinsic relaxation dynamics for a system without external driving in the absence of any collective effects (without agent-agent interactions, only subject to the homing force, the speed controller, and the sigmoid force wrapper described in Sect. II.1), the time step used is $\Delta t=0.02$. The system relaxes from an initial, uniformly random distribution to a steady state, analogous to a damped harmonic oscillator. (a) Structural relaxation of a population of non-interacting agents measured as the mean radial distance to the center of the simulation box, normalized by the initial radial distance at time $t_{0}=0$. Sustained oscillations around the center of the simulation box (underdamped regime, cross symbol) become suppressed when moving towards higher speed-controller strengths $K_{sc}$ and lower target agent speeds $s$ (diamond symbol). Critical damping is situated between the square and triangle symbols, at the threshold where a single oscillation disappears during the initial phase of the relaxation. (b) Corresponding dynamical relaxation: the relative deviation of agent speed from the set target agent speed $s$. Error bands indicate the standard error of the mean over $N=200$ agent trajectories. For reference, the Lyapunov timescale of the Lorenz-63 driver is shown in the time plots (a)-(c), i.e., $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [100] as a gray dashed line. (c,d) Dynamical correlations under driving for driven agents subject to the full active matter model described in Sect. II.1. (c) The absolute value of the velocity auto-correlations for selected damping dynamics. The optimal dynamics (around triangle) entail unique features described in the main text. (d) The connected velocity correlation function (CVC) for different parameter combinations. The optimally damped regime shows the strongest connected velocity fluctuation correlations and anti-correlations, measured spatially via a radial distance around each particle.
@@ III.2.1 Intrinsic damping dynamics: No driving
Fig. 3b shows the intrinsic microscopic relaxation dynamics as the relative deviation of agent speed from the set target agent speed from the initial condition. Seen in both (a) and (b), critical damping is situated between the square and triangle symbols, at the threshold where a single oscillation disappears during the initial phase of the relaxation, which becomes exponential for both the structure and dynamics at the square symbol. Further, the microscopic relaxation dynamics around the diamond symbol has an appearance of a power-law regime with superimposed oscillations. The oscillations decay and the power-law duration drastically shortens when introducing particle interactions Fig. 20d). In (b), the power-law(-like) regime is retained for the square and triangle symbols (nearly critically damped), and disappears for the circle symbol (overdamped). Small oscillations persist at later phases at the steady state. We refer to the crossover case near critical damping (triangle symbol) as the “optimal damping” regime for active matter RC Fig. 2; it is the first overdamped parameter combination below critical damping.
For the optimal damping case (triangle symbol), there is a rapid relaxation to a low value in the structure (a), while in the dynamics, there is an intermediate stage exhibited that appears power-law-like. Intriguingly, for both cases of triangle and circle (more overdamped), the steady state values in dynamics are reached at the same time and with the same value. Yet, this is an effect of the lack of particle interactions: In the case of interacting particles, shown in the Supplementary Fig. 20d,f, the optimally damped system (triangle) indeed converges earlier and more efficiently in dynamics than the case of the circle symbol. The combination of intrinsic exponential decay and a power-law-like regime, hence, seems to have some utility for relaxing the dynamics (velocities) more effectively. It could be a dynamical feature enabling favorable properties of the system, discussed later in Sect. IV. Refer to Tab. 6 for the corresponding videos for the interacting agents case.
@@ III.2.2 Correlative dynamics under driving
In Fig. 3c, the absolute value of the velocity auto-correlations is plotted for selected variations of the speed-controller parameters (symbols) as shown in Fig. 2a. (Here, oscillations are turned into positive peaks with steep zero-crossings on the long-scale; the first peak always corresponds to a negative correlation, and positive peaks alternate thereafter.) The optimal dynamics (around the triangle symbol) entail initial short-time viscous or overdamped motion (an initial exponential decay). This is followed by harmonic-like signals with rounded peaks, exhibiting a shorter frequency closely matching that of the external driving force (the time of the first zero crossing of the VAC). Intriguingly, a near-coincidence of the driver’s Lyapunov timescale, $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [100], with the minimum of the second (negative) peak.
@@
Moving towards higher damping strengths, we observe much shorter timescales between extrema or zero crossings (i.e., shorter oscillations in the VAC). The optimally damped regime (triangle symbol), for example, displays an early first minimum and overall three minima in the absolute VAC within the Lyapunov time of the chaotic driver, compared to only two minima in the underdamped regime. Moreover, the peaks are nearly paraboloid-shaped for the former. Overall, around optimal damping, a diversity of discernible types of dynamics appears (i.e., short-time exponential relaxation and harmonic-like oscillations). Note here that at the crossover (green circle) from the critically damped to the arrested regime, there are additional high-frequency oscillations in the VAC due to its vicinity to the arrested regime (nabla symbol, light green), where particles undergo arrested, back-and-forth oscillatory motion.
@@
Another important quantity that expresses spatial correlations of dynamical fluctuations within the agent system is the CVC (Eq. (19)) [95]. It measures how strongly a change in the velocity of one agent correlates with a change in the velocity of neighboring agents at a relative radial distance $r$. Fig. 3d shows that the critically damped regime (between square and triangle symbols) is characterized by the strongest velocity correlations with next neighbors and the strongest anti-correlations at a radial distance of about $r\approx 3.5$. We also note the short correlation length $r_{0}$ (the first zero crossing of the CVC). A stronger initial decay observed for optimal damping means that fluctuations in velocity rapidly turn from positive into negative correlations.
@@ III.3 Emergent mechanisms in the optimal regime
Figure 4: Phenomenology of optimally damped soft matter systems under driving. Displayed are the response dynamics of $N=500$ agents under quasi-stationary and abrupt driving, which are both part of Lorenz-63 dynamics. Upper panel: Under adiabatic, quasi-stationary driving, agents form a stable interface around the driver, established by the global attractive force to the center of the simulation box (homing force) and the local repulsive force from the driver. Lower panel: Abrupt driving causes the agent-driver interface to break. The driver dashes through the swarm, causing local shear thinning. A new interface forms once the driver slows down. Self-healing through viscous backflow converts the former interface to bulk. The driver tail length corresponds to five integration time steps ($0.15\Delta t=$ 0.1), which is also the time between each snapshot. The snapshots were taken at times $t_{i}=\{1.0,\dots,1.4\}$ (quasi-stationary response) and $t_{i}=\{3.0,\dots,3.4\}$ (highly non-linear response); the Lorenz-63 driving protocol is the same as shown in Fig. 2c. Simulation parameters correspond to those marked by the triangle symbol in Fig. 2a ($K_{sc}=0.02069$, $s=0.04833$). Refer to Supplementary Video Tab. 2/1 for the corresponding video.
@@ III.4 Correlative spatiotemporal velocity fluctuations indicate high performance regimes
Figure 5: The near-critically damped regime shows strong correlations and anti-correlations of agent velocity fluctuations.. (a) The dynamical susceptibility $\chi$ (Eq. 21) captures the intensity and spatial reach between velocity fluctuation correlations (Eq. 16, deviations of individual agent velocities from the system’s center of mass velocity). The high predictive performance regime found in Fig. 2a correlates with the $\chi>5.0$ region. (b) The anti-correlative dynamical response $\text{CVC}(r_{\text{min}})$, the minimum of the connected velocity correlation function (Eq. 19) shown in Fig. 3d), quantifies the maximum extent of anti-correlations between agent velocity fluctuations and correlates with the predictive performance Fig. 2a. The strongest signal is observed just below critical damping at the triangle symbol. Extrema around the “L” symbol could stem from strong rotational modes that were not considered for velocity fluctuation computations. (c) Visualizations of velocity fluctuations for different parameter combinations (symbols) in the speed-controller parameter scans of (a,b). Arrow length indicates strength, and color indicates orientation of velocity fluctuations. Refer to Tab. 4 for corresponding videos.
@@
A different viewpoint is considering the strength of anti-correlations at the first minimum (at $r_{min}$) of the CVC in Fig. 5b. We observe that at the optimal parameter combination (triangle symbol), the greatest degree of anti-correlative strength is reached. It offers a complementary indicator for high predictive performance compared to our calculated dynamical susceptibility, particularly at the crossover region between underdamped/ overdamped and optimally damped regimes (the yellow dashed diagonals), respectively. At the upper-right-hand corner of the diagram, we observe extrema in $\chi$ and $\text{CVC}(r_{\text{min}})$, around the “L”/hexagon symbol. Strong rotational modes (see Supplementary Fig. 15b) that occur in these regimes and that were not accounted for when computing velocity fluctuations could explain these extrema.
Examining the velocity fluctuation snapshots in Fig. 5c shows that in the underdamped regime (cross symbol), the amount of correlation among velocity fluctuations – in their spatial distribution – is low. As the speed-controlling force increases (diamond symbol), the velocity fluctuations become more correlated, indicated by a smoother color gradient of the arrows – a more gradual change of orientation of neighboring velocity fluctuations. Only near the optimal and overdamped cases, represented by the triangle and circle symbols, do we visually recognize the localization of prominent anti-correlated velocity fluctuations. Here, those agents that are close to the source of the disturbance, the driver, experience strong velocity fluctuations. The remainder of the swarm features velocity fluctuations on a much smaller scale, creating a neutral background for a sharp contrast of velocity fluctuations localized near the driver. In contrast, in the underdamped regime (cross symbol), driver-induced velocity fluctuations cannot be distinguished as easily from intrinsic velocity fluctuations of the system. This highlights the favorability of the following duality for active matter reservoir computing: a well-defined swarm “ground state” and local, driver-induced agent excitations. This finding solidifies Lymburn et al.’s observation that there must be a functional dependency of the active matter response on the input signal [62] – which is here stronger in the critically damped regime compared to the underdamped regime.
@@ III.7 Revisiting the role of particle alignment
Figure 8: Alignment forces can improve performance in the speed-controller regime of Lymburn et al. (2021), but not in the optimally damped regime. (a,b) Scans over the parameters of the agent-agent alignment interaction – the interaction (cut-off) radius $r_{a}\in[10^{-1};10^{1}]$ around each agent and the strength $K_{a}\in[10^{-3};10^{1}]$ of the interaction – using (a) speed-controller settings of Lymburn et al. (2021) ($K_{sc}=2.0$, $s=10.0$) [62] and (b) critically damped speed-controller settings ($K_{sc}=0.02069$, $s=0.04833$; triangle symbol in Fig. 2a). (a) Global, weak alignment leads to the formation of droplets, which increases the coherence of the active matter system and its predictive performance. High alignment strengths induce local (intermediate $r_{a}$) and global (high $K_{a}$) frustration in the system. The hexagon/“L” symbol corresponds to the parameter combination used in Ref. [62] ($K_{a}=0.01$, $r_{a}=1.0$; see Supplementary Video Tab. 1/2 for the corresponding snapshot and video). (b) In the critically damped regime, the presence of an alignment force with a strength $K_{a}\gtrapprox 0.1$ decreases predictive performance in dependence on the reach of the interaction. This demonstrates that alignment is detrimental to predictive performance in an optimally damped setting. (c) Snapshots of driven swarms at different parameter combinations indicated by symbols in sub-figure (a). Refer to Tab. 12 for corresponding videos.
@@
Higher interaction radii $r_{a}$ and alignment strengths $K_{a}$ then lead to the formation of an active crystal droplet phase (triangle symbol). This phase yields performances of $P=0.798$ and shows high polar order (see Supplementary Fig. 29a). From Supplementary Video Tab. 12/1 we observe that the long-ranging weak alignment force creates a condensed swarm. Here, the alignment force effectively suppresses driver-induced agent inertia, and repelled agents quickly return to the flock. This leads to the formation of an effectively critically damped droplet with more consistent dynamics upon driving, which is a desirable property for reservoir computing with soft matter as analyzed in Sect. III.1. One reason why the performance could be lower here compared with the actual critically damped regime (shown in Fig. 2a) is that the droplet is not confined to the center of the simulation box. Due to the high mean agent speed, a result of the chosen speed controller settings, the droplet rotates around the box center as an active crystal. This could reduce the prediction quality because similar driving patterns could lead to different swarm states due to this inherent rotational swarm motion. Furthermore, we note an interesting property of the flock: for slow driving, we observe a rotating active crystal that avoids the driver, but is only weakly perturbed by it. For abrupt driving, the active crystal flock aligns its direction of motion to that of the driver. This additional information processing mechanism could improve the predictive quality of the droplet because information about recent driving is conserved in the swarm.
@@
To investigate the importance of the alignment force in the critically damped regime that we presented in Sect. III.1, we perform another parameter scan in Fig. 8b using the previously found optimal (near-critically damped) speed-controller settings ($K_{sc}=0.02069$, $s=0.04833$; triangle symbol in Fig. 2a). The performance heatmap reveals that weak to intermediate alignment does not improve predictive performance in the critically damped regime. We report a maximum performance of $P=0.884$ here, which is equal to the highest performance measured in the speed-controller scan in Fig. 2a. For high alignment strengths and intermediate to high alignment radii, the performance decreases down to zero, as observed before. Overall, this supports the hypothesis that near-critically damped dynamics are optimal because alignment interactions cannot improve the quality of response dynamics and thus only deteriorate the predictive quality. Importantly, this suggests that we can remove alignment as an ingredient in the quest to find the highest-performance soft matter reservoirs.

### v3

@@ III.2 Signatures from microscopic dynamics with and without driving
Figure 3: Characterizing intrinsic microscopic relaxation dynamics (a,b) as well as dynamical correlations under driving (c,d). Symbols correspond to parameter combinations changing speed-controller settings $(K_{sc},s)$ (and resulting damping dynamics) as shown in Fig. 2. (a,b) Intrinsic relaxation dynamics for a system without external driving in the absence of any collective effects (without agent-agent interactions, only subject to the homing force, the speed controller, and the sigmoid force wrapper described in Sect. II.1), the time step used is $\Delta t=0.02$. The system relaxes from an initial, uniformly random distribution to a steady state, analogous to a damped harmonic oscillator. (a) Structural relaxation of a population of non-interacting agents measured as the mean radial distance to the center of the simulation box, normalized by the initial radial distance at time $t_{0}=0$. Sustained oscillations around the center of the simulation box (underdamped regime, cross symbol) become suppressed when moving towards higher speed-controller strengths $K_{sc}$ and lower target agent speeds $s$ (diamond symbol). Critical damping is situated between the square and pyramid symbols, at the threshold where a single oscillation disappears during the initial phase of the relaxation. (b) Corresponding dynamical relaxation: the relative deviation of agent speed from the set target agent speed $s$. Error bands indicate the standard error of the mean over $N=200$ agent trajectories. For reference, the Lyapunov timescale of the Lorenz-63 driver is shown in the time plots (a)-(c), i.e., $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [110] as a gray dashed line. (c,d) Dynamical correlations under driving for driven agents subject to the full active matter model described in Sect. II.1. (c) The absolute value of the velocity auto-correlations for selected damping dynamics. The optimal dynamics (around pyramid) entail unique features described in the main text. (d) The connected velocity correlation function (CVC) for different parameter combinations. The optimally damped regime shows the strongest connected velocity fluctuation correlations and anti-correlations, measured spatially via a radial distance around each particle.
@@ III.2.1 Intrinsic damping dynamics: No driving
Fig. 3b shows the intrinsic microscopic relaxation dynamics as the relative deviation of agent speed from the set target agent speed from the initial condition. Seen in both (a) and (b), critical damping is situated between the square and pyramid symbols, at the threshold where a single oscillation disappears during the initial phase of the relaxation, which becomes exponential for both the structure and dynamics at the square symbol. Further, the microscopic relaxation dynamics around the diamond symbol has an appearance of a power-law regime with superimposed oscillations. The oscillations decay and the power-law duration drastically shortens when introducing particle interactions Fig. 20d). In (b), the power-law(-like) regime is retained for the square and pyramid symbols (nearly critically damped), and disappears for the circle symbol (overdamped). Small oscillations persist at later phases at the steady state. We refer to the crossover case near critical damping (pyramid symbol) as the “optimal damping” regime for active matter RC Fig. 2; it is the first overdamped parameter combination below critical damping.
For the optimal damping case (pyramid symbol), there is a rapid relaxation to a low value in the structure (a), while in the dynamics, there is an intermediate stage exhibited that appears power-law-like. Intriguingly, for both cases of pyramid and circle (more overdamped), the steady state values in dynamics are reached at the same time and with the same value. Yet, this is an effect of the lack of particle interactions: In the case of interacting particles, shown in the Supplementary Fig. 20d,f, the optimally damped system (pyramid) indeed converges earlier and more efficiently in dynamics than the case of the circle symbol. The combination of intrinsic exponential decay and a power-law-like regime, hence, seems to have some utility for relaxing the dynamics (velocities) more effectively. It could be a dynamical feature enabling favorable properties of the system, discussed later in Sect. IV. Refer to Tab. 6 for the corresponding videos for the interacting agents case.
@@ III.2.2 Correlative dynamics under driving
In Fig. 3c, the absolute value of the velocity auto-correlations is plotted for selected variations of the speed-controller parameters (symbols) as shown in Fig. 2a. (Here, oscillations are turned into positive peaks with steep zero-crossings on the long-scale; the first peak always corresponds to a negative correlation, and positive peaks alternate thereafter.) The optimal dynamics (around the pyramid symbol) entail initial short-time viscous or overdamped motion (an initial exponential decay). This is followed by harmonic-like signals with rounded peaks, exhibiting a shorter frequency closely matching that of the external driving force (the time of the first zero crossing of the VAC). Intriguingly, a near-coincidence of the driver’s Lyapunov timescale, $t^{\text{L63}}_{\text{lyap}}\approx 1.1$ [110], with the minimum of the second (negative) peak.
@@
Moving towards higher damping strengths, we observe much shorter timescales between extrema or zero crossings (i.e., shorter oscillations in the VAC). The optimally damped regime (pyramid symbol), for example, displays an early first minimum and overall three minima in the absolute VAC within the Lyapunov time of the chaotic driver, compared to only two minima in the underdamped regime. Moreover, the peaks are nearly paraboloid-shaped for the former. Overall, around optimal damping, a diversity of discernible types of dynamics appears (i.e., short-time exponential relaxation and harmonic-like oscillations). Note here that at the crossover (green circle) from the critically damped to the arrested regime, there are additional high-frequency oscillations in the VAC due to its vicinity to the arrested regime (nabla symbol, light green), where particles undergo arrested, back-and-forth oscillatory motion.
@@
Another important quantity that expresses spatial correlations of dynamical fluctuations within the agent system is the CVC (Eq. (19)) [105]. It measures how strongly a change in the velocity of one agent correlates with a change in the velocity of neighboring agents at a relative radial distance $r$. Fig. 3d shows that the critically damped regime (between square and pyramid symbols) is characterized by the strongest velocity correlations with next neighbors and the strongest anti-correlations at a radial distance of about $r\approx 3.5$. We also note the short correlation length $r_{0}$ (the first zero crossing of the CVC). A stronger initial decay observed for optimal damping means that fluctuations in velocity rapidly turn from positive into negative correlations.
@@ III.3 Emergent mechanisms in the optimal regime
Figure 4: Phenomenology of optimally damped soft matter systems under driving. Displayed are the response dynamics of $N=500$ agents under quasi-stationary and abrupt driving, which are both part of Lorenz-63 dynamics. Upper panel: Under adiabatic, quasi-stationary driving, agents form a stable interface around the driver, established by the global attractive force to the center of the simulation box (homing force) and the local repulsive force from the driver. Lower panel: Abrupt driving causes the agent-driver interface to break. The driver dashes through the swarm, causing local shear thinning. A new interface forms once the driver slows down. Self-healing through viscous backflow converts the former interface to bulk. The driver tail length corresponds to five integration time steps ($5\Delta t=0.1$), which is also the time between each snapshot. The snapshots were taken at times $t_{i}=\{1.0,\dots,1.4\}$ (quasi-stationary response) and $t_{i}=\{3.0,\dots,3.4\}$ (highly non-linear response); the Lorenz-63 driving protocol is the same as shown in Fig. 2c. Simulation parameters correspond to those marked by the pyramid symbol in Fig. 2a ($K_{sc}=0.02069$, $s=0.04833$). Refer to Supplementary Video Tab. 2/1 for the corresponding video.
@@ III.4 Correlative spatiotemporal velocity fluctuations indicate high performance regimes
Figure 5: The near-critically damped regime shows strong correlations and anti-correlations of agent velocity fluctuations.. (a) The dynamical susceptibility $\chi$ (Eq. 21) captures the intensity and spatial reach between velocity fluctuation correlations (Eq. 16, deviations of individual agent velocities from the system’s center of mass velocity). The high predictive performance regime found in Fig. 2a correlates with the $\chi>5.0$ region. (b) The anti-correlative dynamical response $\text{CVC}(r_{\text{min}})$, the minimum of the connected velocity correlation function (Eq. 19) shown in Fig. 3d), quantifies the maximum extent of anti-correlations between agent velocity fluctuations and correlates with the predictive performance Fig. 2a. The strongest signal is observed just below critical damping at the pyramid symbol. Extrema around the “L” symbol could stem from strong rotational modes that were not considered for velocity fluctuation computations. (c) Visualizations of velocity fluctuations for different parameter combinations (symbols) in the speed-controller parameter scans of (a,b). Arrow length indicates strength, and color indicates orientation of velocity fluctuations. Refer to Tab. 4 for corresponding videos.
@@
A different viewpoint is considering the strength of anti-correlations at the first minimum (at $r_{min}$) of the CVC in Fig. 5b. We observe that at the optimal parameter combination (pyramid symbol), the greatest degree of anti-correlative strength is reached. It offers a complementary indicator for high predictive performance compared to our calculated dynamical susceptibility, particularly at the crossover region between underdamped/ overdamped and optimally damped regimes (the yellow dashed diagonals), respectively. At the upper-right-hand corner of the diagram, we observe extrema in $\chi$ and $\text{CVC}(r_{\text{min}})$, around the “L”/hexagon symbol. Strong rotational modes (see Supplementary Fig. 15b) that occur in these regimes and that were not accounted for when computing velocity fluctuations could explain these extrema.
Examining the velocity fluctuation snapshots in Fig. 5c shows that in the underdamped regime (cross symbol), the amount of correlation among velocity fluctuations – in their spatial distribution – is low. As the speed-controlling force increases (diamond symbol), the velocity fluctuations become more correlated, indicated by a smoother color gradient of the arrows – a more gradual change of orientation of neighboring velocity fluctuations. Only near the optimal and overdamped cases, represented by the pyramid and circle symbols, do we visually recognize the localization of prominent anti-correlated velocity fluctuations. Here, those agents that are close to the source of the disturbance, the driver, experience strong velocity fluctuations. The remainder of the swarm features velocity fluctuations on a much smaller scale, creating a neutral background for a sharp contrast of velocity fluctuations localized near the driver. In contrast, in the underdamped regime (cross symbol), driver-induced velocity fluctuations cannot be distinguished as easily from intrinsic velocity fluctuations of the system. This highlights the favorability of the following duality for active matter reservoir computing: a well-defined swarm “ground state” and local, driver-induced agent excitations. This finding solidifies Lymburn et al.’s observation that there must be a functional dependency of the active matter response on the input signal [72] – which is here stronger in the critically damped regime compared to the underdamped regime.
@@ III.7 Revisiting the role of particle alignment
Figure 8: Alignment forces can improve performance in the speed-controller regime of Lymburn et al. (2021), but not in the optimally damped regime. (a,b) Scans over the parameters of the agent-agent alignment interaction – the interaction (cut-off) radius $r_{a}\in[10^{-1};10^{1}]$ around each agent and the strength $K_{a}\in[10^{-3};10^{1}]$ of the interaction – using (a) speed-controller settings of Lymburn et al. (2021) ($K_{sc}=2.0$, $s=10.0$) [72] and (b) critically damped speed-controller settings ($K_{sc}=0.02069$, $s=0.04833$; pyramid symbol in Fig. 2a). (a) Global, weak alignment leads to the formation of droplets, which increases the coherence of the active matter system and its predictive performance. High alignment strengths induce local (intermediate $r_{a}$) and global (high $K_{a}$) frustration in the system. The hexagon/“L” symbol corresponds to the parameter combination used in Ref. [72] ($K_{a}=0.01$, $r_{a}=1.0$; see Supplementary Video Tab. 1/2 for the corresponding snapshot and video). (b) In the critically damped regime, the presence of an alignment force with a strength $K_{a}\gtrapprox 0.1$ decreases predictive performance in dependence on the reach of the interaction. This demonstrates that alignment is detrimental to predictive performance in an optimally damped setting. (c) Snapshots of driven swarms at different parameter combinations indicated by symbols in sub-figure (a). Refer to Tab. 12 for corresponding videos.
@@
Higher interaction radii $r_{a}$ and alignment strengths $K_{a}$ then lead to the formation of an active crystal droplet phase (pyramid symbol). This phase yields performances of $P=0.798$ and shows high polar order (see Supplementary Fig. 29a). From Supplementary Video Tab. 12/1 we observe that the long-ranging weak alignment force creates a condensed swarm. Here, the alignment force effectively suppresses driver-induced agent inertia, and repelled agents quickly return to the flock. This leads to the formation of an effectively critically damped droplet with more consistent dynamics upon driving, which is a desirable property for reservoir computing with soft matter as analyzed in Sect. III.1. One reason why the performance could be lower here compared with the actual critically damped regime (shown in Fig. 2a) is that the droplet is not confined to the center of the simulation box. Due to the high mean agent speed, a result of the chosen speed controller settings, the droplet rotates around the box center as an active crystal. This could reduce the prediction quality because similar driving patterns could lead to different swarm states due to this inherent rotational swarm motion. Furthermore, we note an interesting property of the flock: for slow driving, we observe a rotating active crystal that avoids the driver, but is only weakly perturbed by it. For abrupt driving, the active crystal flock aligns its direction of motion to that of the driver. This additional information processing mechanism could improve the predictive quality of the droplet because information about recent driving is conserved in the swarm.
@@
To investigate the importance of the alignment force in the critically damped regime that we presented in Sect. III.1, we perform another parameter scan in Fig. 8b using the previously found optimal (near-critically damped) speed-controller settings ($K_{sc}=0.02069$, $s=0.04833$; pyramid symbol in Fig. 2a). The performance heatmap reveals that weak to intermediate alignment does not improve predictive performance in the critically damped regime. We report a maximum performance of $P=0.884$ here, which is equal to the highest performance measured in the speed-controller scan in Fig. 2a. For high alignment strengths and intermediate to high alignment radii, the performance decreases down to zero, as observed before. Overall, this supports the hypothesis that near-critically damped dynamics are optimal because alignment interactions cannot improve the quality of response dynamics and thus only deteriorate the predictive quality. Importantly, this suggests that we can remove alignment as an ingredient in the quest to find the highest-performance soft matter reservoirs.